OpenStudy (anonymous):

the data shows that the bone lengths have a mean of 45.9 cm and standard deviation of 4.2 cm, the overall heights have a mean of 172.7 cm and standard deviation of 8.14 cm, and the correlation between bone length and height is 0.914. what is the slope of the LSRL of height on bone length?

OpenStudy (amistre64):

least squares regression line?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

hmm .... i wonder how we can fit this into the only thing i know for finding the slope of this:\[m=\frac{n\sum xy-\sum x\sum y}{n\sum xx-\sum x\sum x }\]

OpenStudy (amistre64):

i know this is an analysis of variance .... type setup

OpenStudy (anonymous):

I don't know how to put it in that equation

OpenStudy (amistre64):

me either, at the moment, im pretty sure there is some simplified versions that are in the textbook ... all check back in a few minutes

OpenStudy (amistre64):

\[\sigma^2=\frac{\sum(x-\mu)^2}{n}=\frac{\sum x^2-2\mu\sum x+\mu^2\sum1}{n}\] \[\sigma^2=\frac{\sum x^2}{n}-2\mu\frac{\sum x}{n}+\mu^2\frac{\sum1}{n}\] \[\sigma^2=\frac{\sum x^2}{n}-2\mu\mu+\mu^2\frac{n}{n}\] \[\sigma^2=\frac{\sum x^2}{n}-\mu^2\] \[\sigma^2=\frac{\sum x^2}{n}-\frac{\sum x}{n}\frac{\sum x}{n}\] \[\sigma^2=\frac{\sum x^2-\sum x\sum x}{n}\] which is prolly how we get to the parts in the slope setup

OpenStudy (amistre64):

im confusing r for slope ... but they are very similar

OpenStudy (amistre64):

\[m=r\frac{s_y}{s_x}\] where the s parts are the standard deviations, and r is the linear correlation coeff

OpenStudy (amistre64):

any recollection from your material how to find r with the given information?

OpenStudy (anonymous):

no I don't have a book or anything

OpenStudy (amistre64):

"and the correlation between bone length and height is 0.914" r = 0.914 , simply read right over that :/ m = 0.914(8.14)/4.2 assuming bone is along the x axis ...